OpenStudy (anonymous):

Use synthetic substitution to evaluate P(x) = x3 + 3x2 - 6 for x = 5. P(5) =

OpenStudy (ash2326):

airricuhh why don't you try this one? if you get stuck I'll help you

OpenStudy (anonymous):

so you start at 5 3 3 6 and then i don't know what comes under that..

OpenStudy (anonymous):

goes"

OpenStudy (ash2326):

Let me teach you,

OpenStudy (anonymous):

\[5^3+3(5^2)-6=194\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no i don't saljudieh,

OpenStudy (ash2326):

we have x=5, it means we are dividing by x-5 to get the value of P(5) write the coefficients of P(x)= x^3+3x^2-6 coefficients are 1 3 -6

OpenStudy (ash2326):

got till here?

OpenStudy (anonymous):

where did you get the 1 coefficient from? the x right?

OpenStudy (ash2326):

yeah x^3 = 1 3x^2= 3 -6 as it is

OpenStudy (anonymous):

okay, so where do you get the numbers that go under neath,

OpenStudy (ash2326):

now we are using synthetic division first step write this way _______________ 5 | 1 3 -6

OpenStudy (anonymous):

wait ash, i messed up the problem was Use synthetic substitution to evaluate P(x) = x3 + 3x2 - 6 for x = -1. P(-1) =

OpenStudy (ash2326):

now we have _______________ 5 | 1 3 -6 bring the first term 1 as it is down , _______________ 5 | 1 3 -6 1 next multiply , the term brought down by 5 and ____________________________ 5 | 1 3 -6 5(5*1) 1

OpenStudy (anonymous):

:/ im sorry, so it would be _______________ -1 | 1 3 -6

OpenStudy (ash2326):

no we have to evaluate at x=5 so it's _______________ 5 | 1 3 -6

OpenStudy (anonymous):

i know but i typo'd the problem its P(-1) :/

OpenStudy (anonymous):

so it would be 3 5 which is 15 so then would i divide 15 by -6 and get the answer?

OpenStudy (anonymous):

hey arricuhh u wanna learn? watch thishttp: http://www.youtube.com/watch?v=-aTMzg4syy8

OpenStudy (anonymous):

well thanks.

OpenStudy (anonymous):

did it help?

OpenStudy (anonymous):

so is the answer -8 ?

OpenStudy (anonymous):

nope, it did help. but i still got the wrong answer.

OpenStudy (anonymous):

wait is it -5?

OpenStudy (anonymous):

 ok look at this: